In preparing an infusion of the drug trastuzumab, a pharmacist adds 20 ml of bacteriostatic water for injection to a vial containing 440 mg of the mAb. The resulting solution contains trastuzumab, 21 mg/ml. Using the information: Trastuzumab loading dose: 4 mg/kg with 90 min infusion

(a) Calculate the milliliters of solution required for the loading dose for a 165 lb. Patient.

(b) The pharmacist then transfers the calculated volume into an infusion bag containing 250 ml of sodium chloride injection. Calculate the quantity of trastuzumab administered in mg/min.

(c) Calculate the drip rate in drops/min, using an administered set that delivers 20 drops/ml
Answer #1

Drug Trastuzumab 440mg. Here pharmacist adds 20 ml bacteriostatic water into it and resulting 21mg/ml of the drug.

hence;

WHAT WE HAVE = 21mg/ml

Loading dose = 4mg/Kg

Weight of the patient = 165lbs

1lbs = 0.454Kg

165lbs = 0.454 × 165 =74.9Kg

Dose to be given for patient according to loading dose = 4 × 74.9 = 299.6mg

WHAT WE WANT = 299.6mg

a) millilitres of solution to be administered = WHAT WE WANT/ WHAT WE HAVE

Millilitres of solution to be administered

= 299.6mg/21mg/ml. = 14.3ml